Superior Micro Products uses the weighted average method in its process costing system. During January, the Delta Assembly DePlease answer 1-4 :)
Answer #1

1. Equivalent Units = Ending WIP Inventory * Completion %

Equivalent Units for materials = 2600 * 70% = 1,820 Units

Equivalent Units for Labor = 2600 * 50% = 1300 Units

Equivalent Units for Overhead = 2600 * 50% = 1300 Units

2. Computation of cost of ending Inventoy in WIP

cost of ending inventory for materials = 1820 x 12.50 = $22,750
cost of ending inventory for labor = 1,300 x 4.80 = $6,240
cost of ending inventory for overhead = 1,300 x 8.0 = $10,400

Total cost of ending inventory = $22,750 + $6,240 + $10,400= $39,390

3 . Cost of Units transferred to next department

cost of ending work in process inventory for materials = 26,100 x 12.50 = $326,250
cost of ending work in process inventory for labor = 26,100 x 4.80 = $125,280
cost of ending work in process inventory for overhead = 26,100 x 8.0 = $208,800

Total cost of ending work in process inventory = $326,250 + $125,280 + $208,800 = $660,330

4. Cost reconciliation for january

Total cost to be accounted for $699,720 = cost of ending work in process inventory $39,390 + cost of units completed and transferred out $660,330

Total cost accounted for $699,720 ( total cost to be accounted for = cost of beginning work in process inventory + cost added to production during the period) ( Given in the question thus the same is reconcilied)